OpenStudy (anonymous):

Solve the system of equations using elemination 2x-3y+z=0 2x+y+z=12 y-z= 4

OpenStudy (anonymous):

What have you done so far? Do you understand what it's asking you to do?

OpenStudy (anonymous):

Not really I've done some problem with 2 equations but not three

OpenStudy (anonymous):

It's the same concept but with one more step. You can begin anywhere, but I'll be using equation 3 as it's easiest to rearrange to make a variable the subject. \[y=4+z\] We can use this to eliminate y from the other two equations, leaving us with two equations in terms of x and z. Equation 1: \[0=2x-3(4+z)+z=2x-12-3z+z=2x-2z-12\]\[2x-2z=12\]\[x-z=6\] Equation 2: \[12=2x+y+z=2x+(4+z)+z=2x+4+2z\]\[2x+2z=8\]\[x+z=4\] Solve these two equations for x and z, then use y=4+z to get y.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

My pleasure. If you want to check your answers, I believe it's x=5, y=3, z=-1.